# CATD: Unified Representation Learning for EEG-to-fMRI Cross-Modal Generation

### v1

@@ 3.1.1 Dataset
The experiments were performed on the following datasets: the motor imagery dataset[36], the NODDI dataset[37, 38], and the EEG dataset of Parkinsonian patients[39]. For evaluating the reconstruction results, the XP1 part of the motor imagery dataset was used, which contained data from 10 subjects (2 females and 8 males, mean age: 28.4 years $\pm$ 10.6 years). Their paired 64-channel EEG and whole-brain fMRI scans were acquired simultaneously using a block design, with each block consisting of a 20-second rest and a 20-second motor imagery. The motor imagery task contained 48 blocks per subject during the test duration. The NODDI dataset included simultaneous resting-state 64-channel EEG and whole-brain fMRI scan data from 17 adult volunteers (11 males and 6 females, mean age: 32.84 $\pm$ 8.13 years). The EEG dataset of the Parkinson’s patients contained 64-channel resting-state brain EEG data from 8 subjects (4 males and 4 females, mean age: 74.25 years $\pm$ 8.75 years). EEG data in all datasets were obtained using the international 10-20 lead system.
@@ 3.1.2 Implementation Detail
The experiments were conducted on a server platform equipped with two Nvidia Tesla A800 compute cards. For model parameters, the depth of the Transformer Block was set to 12, the hidden space dimension of the patch was 768, and the number of heads in the attention network was 12. Training was performed using the Adam optimizer with an initial learning rate of 0.0001, a batch size of 8, and over 1000 epochs. To compute the experimental classification metrics, a five-fold cross-validation method was employed to ensure the reliability of the results. For calculating other quantitative metrics, five independent experiments were conducted, and the results were averaged to ensure data accuracy and stability.
@@ 3.1.3 Metrics
A variety of categorical metrics were employed, including Accuracy (ACC), Precision (PRE), Sensitivity (SEN), and F1-score, to demonstrate the performance of the synthesized results in downstream tasks. Additionally, to evaluate reconstruction quality in the spatial dimension, Root Mean Square Error (RMSE) and Structural Similarity Index (SSIM) were used. For the temporal dimension, Cosine Similarity and Concordance Correlation Coefficient (CCC) were used to assess reconstruction quality. For the temporal super-resolution experiments, Signal-to-Noise Ratio (SNR) was also used to evaluate the effectiveness of the synthesized signal, with SNR values being base-10 logarithms. To present the results more intuitively, graphical representations of classification and reconstruction performance in different experiments were provided. For example, in the medical decision support experiment, the potential application of the method in medical tasks was illustrated through difference maps of BOLD function maps.
3.2 Evaluation of the reconstructed BOLD signal
@@
To assess the effectiveness of the EEG-to-BOLD signal reconstruction, both synthetic and real BOLD signals were utilized to differentiate between subjects’ performance in motor imagery and resting states. The brain activity of the subjects was divided into 20-second chunks for both resting and motor imagery tasks. A clear distinction between these states by the model indicated effective learning of brain activity patterns. As shown in Fig.3(a), the cross-modal synthetic BOLD signals outperformed the real data in accuracy and precision. When combined with real data, the classification metrics improved significantly, although recall remained similar to the real data. Ablation experiments demonstrated that CAB enhances the model’s ability to use EEG features to constrain BOLD signal synthesis, improving classification metrics compared to models without the CAB.
@@
Spatial (RMSE, SSIM) and temporal (cosine similarity, CCC) metrics of the synthesized BOLD signal were also calculated for the three motion picture states. The results in Fig.4 show low RMSEs (all below 0.1) and high structural similarity indices (all above 0.6735). The cosine similarity is approximately 0.85, and the CCC value is about 0.8, indicating a high spatio-temporal correlation with the real signal. The ablation experiments (e.g., Fig.3(a) and the No CAB section in Fig.4) confirm the effectiveness of the CAB in improving signal quality.
@@
Figure 4: Results of quantitative spatial and temporal metrics for the synthetic BOLD signal in three states of motor imagery. The upper half shows the metrics obtained using the CATD framework we propose, and the lower half shows the metrics obtained after ablation experiments on conditioned blocks.
@@
To further illustrate the capability of the proposed model in learning high-dimensional heterogeneous brain activity features and achieving cross-modal transitions, t-SNE plots were used (Fig.5). The distribution of BOLD signals synthesized with the full CATD framework more closely matches the real signal distribution, underscoring the critical role of the CAB in enhancing signal quality.
@@
Figure 5: (a) Shows the tSNE plots of the distribution of the synthesized BOLD signal versus the real signal in the CATD framework for the three motor imagery states. (b) Demonstrates the t-SNE plot of the synthetic BOLD signal versus the real signal distribution after ablation of the conditioned block.
@@ 3.3 Evaluation of temporal resolution enhanced BOLD signals
To verify that the CATD framework can leverage the high temporal resolution of EEG signals to achieve the temporal super-resolution of BOLD signals, temporal resolution enhancement experiments were conducted. The proposed DTFS was used for EEG to achieve triple temporal super-resolution, meaning that the temporal resolution of the reconstructed BOLD signal was three times that of the actual BOLD signal. Since the application scenario for temporal super-resolution typically involves enhancing the existing BOLD signal rather than generating it in the absence of a BOLD signal, the original low temporal resolution BOLD signal was used as a known condition. The enhanced high temporal resolution BOLD signal was obtained by constraining the generated signal using the low-resolution original BOLD signal. The hyper-temporal resolution reconstruction results were similarly evaluated in three different motor imagery states, as shown in Table 1.
@@
Table 1: Results of the time series similarity between the reconstructed triple time-resolved BOLD signal and the real BOLD signal in three motion imagery states.
Cosine Similarity CCC
eegNF 0.98998$\pm$0.00008 0.98589$\pm$0.00003
fmriNF 0.99187$\pm$0.00007 0.98714$\pm$0.00006
eegfmriNF 0.98740$\pm$0.00005 0.98209$\pm$0.00002
Table 2: Comparison of SNR between real BOLD signals with low temporal resolution and synthesized BOLD signals with high temporal resolution in three motion imagery states.
Real Synthetic
eegNF 12.9640 12.9995$\pm$0.0103
fmriNF 13.4383 13.3112$\pm$0.0137
eegfmriNF 12.4715 13.0708$\pm$0.0533
@@
An important advantage of high temporal resolution is the potential for signal-to-noise ratio (SNR) enhancement. A higher SNR implies a higher proportion of useful information in the signal relative to the noise and signals with a higher SNR can more reliably reflect the actual physiological activity situation. In the experiments, the SNR values of real and reconstructed BOLD signals were shown in Table 2. From the results, it could be seen that the signal-to-noise ratios in the two states were improved except for fmriNF, the motion imagery state. This was also consistent with the results demonstrated by t-SNE above, i.e., the results were worse in the fmriNF state and better in the other two states.
@@
Figure 6: Visualization of real BOLD signal points with low time resolution and synthetic BOLD signal curves with high time resolution. The graph shows that the synthetic signal has the same trend as the real signal.
@@
Data visualization was also performed. In the parietal region, where the correlation of motor imagery was strong, a node at the cortex corresponding to the C3 electrode portion of the EEG international standard lead system was selected. Curve graphs were used to display the signal intensities of the reconstructed high temporal-resolution time-series signals and the real signals at this node at the corresponding time points. As shown in Fig.6, the reconstructed high temporal resolution signal and the real signal exhibited a high degree of similarity in trend. This further suggests that the hyper-temporal resolution reconstruction results obtained using the CATD framework can effectively reflect the trend of blood oxygenation.
3.4 Evaluation of the effect of temporal features on the reconstructed signals
@@
Different frequency bands of EEG data reflect various types of brain activity, and these differences can significantly impact the results. The effect of EEG frequency bands on BOLD signal reconstruction results was investigated. The EEG signals of different frequency bands were obtained by band-pass filtering, and BOLD signal reconstruction was performed using these signals. As shown in Fig.3(b), the beta and gamma bands, which are related to motor imagery, performed better in the classification index. Additionally, the timing metrics of the reconstruction results were calculated. As shown in Fig.7, the gamma band performed well in the timing metrics, while the beta band showed lower timing metrics.
@@
The reasons for these results are analysed below. The brain’s state changes in preparation for or during imagined movements are generally slower and smoother, leading to event-related desynchronization (ERD) in the beta band[40]. This phenomenon may explain why the beta band excels in motor imagery and resting state categorization, even surpassing the full band’s categorization performance. However, this characteristic also causes the reconstructed results of the beta band to differ considerably in temporal similarity, resulting in a lower temporal index compared to the blood oxygenation activity reflected by the full-band brain electrical activity, i.e., the real BOLD signal. On the other hand, the gamma band is closely associated with higher cognitive and perceptual functions. In motor imagery tasks, the brain’s requirement to understand instructions and make judgments involves higher cognitive and perceptual functions[41]. Therefore, the gamma band most closely matches the real results in both the categorization index and timing index. The other three low-frequency bands: delta, theta, and alpha, performed poorly on both the classification metric (e.g., Fig.3(b)) and the temporal similarity metric (e.g., Fig.7), suggesting that these bands negatively affect the reconstruction results by containing less information about the correlation between the EEG and the BOLD signal.
@@ 3.5 Support for medical decision-making
To demonstrate the potential application of the proposed CATD framework in the medical domain, cross-modal reconstruction experiments were conducted on Parkinson’s patients to support medical decision-making. In this part of the experiment, the framework was first trained on the NODDI dataset containing paired EEG and BOLD fMRI data from healthy subjects at resting state. Subsequently, another dataset containing only resting-state EEG data from Parkinson’s patients was used to reconstruct the cortical BOLD functional maps of these patients using the trained cross-modal generation model.
Predictions of Parkinson’s disease were performed using the true EEG signal, the reconstructed BOLD signal, and a combination of the true EEG signal and the reconstructed BOLD signal. The classification metrics are shown in Fig.3(c). Results indicate that the simultaneous use of reconstructed BOLD signals and recorded EEG signals significantly improves the accuracy of Parkinson’s disease prediction and related metrics. This suggests that the CATD framework effectively captures the potential connection between BOLD signals and EEG signals and can be applied across different datasets, which is crucial for disease diagnosis. Consequently, the framework not only enhances the accuracy of existing diagnostic methods but also provides new tools and methodologies for the diagnosis and research of other neurological diseases.
In order to fully utilize the advantage of our CATD framework, i.e., to obtain BOLD signals with high spatial resolution without the condition of fMRI detection, we performed a difference analysis of reconstructed BOLD functional maps in two healthy subjects and two Parkinson’s patients. As shown in Fig.8, both showed significant abnormalities in the brain regions marked by red circles. This region is the lingual gyrus, the medial occipito-temporal gyrus and the lateral occipito-temporal gyrus[42], which, in patients with Parkinson’s disease, shows significant structural changes[43], and in patients with PD accompanied by visual hallucinations, the atrophy of these three brain regions correlates with the severity of visual hallucinations[44]. This result demonstrates the potential application of the proposed method in localizing regions of abnormal brain activity and further proves its practical value in medical diagnosis.
@@ 4 Discussion
Cross-modal neuroimage synthesis is becoming crucial in neuroanalysis research. For the first time, we propose the CATD framework based on diffusion models to achieve cross-modal synthesis of temporal functional neuroimages, enabling the conversion of EEG to BOLD signals. This approach addresses limitations in BOLD acquisition, such as the inability of patients with metal implants to undergo fMRI scans. Enhancing BOLD temporal resolution facilitates the study of sub-millimeter cortical structures and activities[45], improves the quality and reliability of functional connectivity and task-driven fMRI research through higher sampling rates[46], and allows for more accurate detection and localization of functional activation regions, capturing transient neural activities[47]. This work provides novel insights for functional neuroimage synthesis[48]. Next, we will analyze the experimental results in detail to illustrate the performance of our framework.
@@
In the experiments on the motor imagery dataset, our CATD framework demonstrates its superior performance in cross-modal feature learning and synthesis by detecting categorical metrics for both motor imagery states and resting states. Specifically, the framework is able to capture the features of EEG and BOLD signals well and perform effective cross-modal synthesis. This improved performance can be attributed to the ability to capture complex modes and integrate different modal features in the CATD framework, which enables it to achieve accurate transitions between various signal features. Ablation experiments further validate the effectiveness of the proposed CAB in aligning high-dimensional mismatched data pairs. CAB can learn and capture valid links between EEG and BOLD signals, ensuring the robustness of the model in cross-modal feature capture and synthesis.
@@
In the temporal super-resolution experiments, the results exhibit a high correlation in all three motion imagery states, while the signal-to-noise ratio is improved in both states compared to the original signal. This suggests that temporal resolution enhancement is indeed feasible and can provide a more detailed and accurate temporal representation of the BOLD signal.
@@
In order to further verify which EEG frequency bands have a greater impact on the results and have the potential to improve the efficiency of cross-modal synthesis, a frequency band analysis was performed. It was found that the beta band had the greatest impact on the classification results, while the gamma band performed the best in terms of similarity to the real signal. This can be explained by the properties of EEG signals. Lower frequency bands such as beta show event-related desynchronization (ERD) phenomenon in motor imagery tasks and therefore perform better in classification metrics. Whereas the higher cognitive functions involved in the motor imagery task are mainly determined by the gamma band, the reconstruction results of the gamma band therefore show the highest similarity to the original signal.
@@
While the results are promising, achieving accurate mapping from EEG to BOLD signals still presents challenges. Fig.9 illustrates a typical failure case, where the lower regions of the reconstructed BOLD map significantly differ from the real image. This discrepancy is likely due to the low spatial resolution and relatively low signal-to-noise ratio (SNR) of EEG signals. Even though our model successfully learns the relationship between EEG and BOLD signals, the process of upsampling to generate high-resolution images introduces inaccuracies. Addressing these issues of spatial resolution and SNR will be crucial for enhancing the precision of cross-modal synthesis in future work.
Figure 9: Typical failure case of our method. It can be seen that regions such as the lower part of the BOLD function map (near the frontal lobe) are more differentiated
Although the limited number of paired data subjects constrained the training, the amount of training data was increased by sampling and pairing the first 6 seconds of each frame of EEG data with a functional map of the BOLD signal through a data segmentation technique. This data augmentation approach positively impacted the model’s performance. All experiments were conducted across different subjects. While the results are not yet perfect, they still indicate a more favorable outcome.

### v3

@@ 3.1.1 Dataset
The experiments were performed on the following datasets: the motor imagery dataset[46], the NODDI dataset[47, 48], and the EEG dataset of Parkinson’s patients[49]. For evaluating the generated results, the XP1 part of the motor imagery dataset was used, which contained data from 10 subjects (2 females and 8 males, mean age: 28.4 years $\pm$ 10.6 years). Their paired 64-channel EEG and whole-brain fMRI scans were acquired simultaneously using a block design, with each block consisting of a 20-second rest and a 20-second motor imagery. The training set included data from 7 subjects (IDs xp101 to xp103, xp107 to xp110), while the test set consisted of data from 3 subjects (IDs xp104 to xp106). This division ensured the independence of the training and testing samples, providing a reliable assessment of the model’s generalization performance across participants. The motor imagery task contained 48 blocks per subject during the test duration. The NODDI dataset included simultaneous resting-state 64-channel EEG and whole-brain fMRI scan data from 17 adult volunteers (11 males and 6 females, mean age: 32.84 $\pm$ 8.13 years). For evaluation, data from 2 participants (IDs 48 and 49) were selected as the test set, while the remaining data were used for training. This division allowed the model to accurately generate BOLD signals from EEG data. The EEG dataset of the Parkinson’s patients contained 64-channel resting-state EEG data from 8 subjects (4 males and 4 females, mean age: 74.25 years $\pm$ 8.75 years). Due to the small sample size and significant clinical variability, data from 4 representative participants were selected based on factors such as age and gender for testing, highlighting the clinical potential of the method in real-world applications. EEG data in all datasets were obtained using the international 10-20 lead system.
@@ 3.1.2 Implementation Detail
The experiments were conducted on a server platform equipped with two NVIDIA Tesla A800 compute cards. For model parameters, the depth of the Transformer Blocks was set to 12, the hidden space dimension of the patch was 768, and the number of heads in the attention network was 12. Training was performed using the Adam optimizer with an initial learning rate of 0.0001, a batch size of 8, and 1000 epochs. To compute the experimental classification metrics, a five-fold cross-validation method was employed to ensure the reliability of the results. For calculating other quantitative metrics, five independent experiments were conducted, and the results were averaged to ensure data accuracy and stability.
@@ 3.1.3 Metrics
Variety of categorical metrics were employed, including Accuracy (ACC), Precision (PRE), Sensitivity (SEN), and F1-score, to demonstrate the performance of the synthesized results in downstream tasks. Additionally, to evaluate generation quality in the spatial dimension, Root Mean Square Error (RMSE) and Structural Similarity Index (SSIM) were used. For the temporal dimension, Cosine Similarity and Concordance Correlation Coefficient (CCC) were used to assess generation quality. For the temporal super-resolution experiments, Signal-to-Noise Ratio (SNR) was also used to evaluate the effectiveness of the synthesized signal, with SNR values being base-10 logarithms. To present the results more intuitively, graphical representations of classification and generation performance in different experiments were provided. For example, in the medical decision support experiment, the potential application of the method in medical tasks was illustrated through difference maps of BOLD function maps.
3.2 Evaluation of the generated BOLD signal
@@
To assess the effectiveness of the EEG-to-BOLD signal generation, both synthetic and real BOLD signals were utilized to differentiate between subjects’ performance in motor imagery and resting states. The brain activity of the subjects was divided into 20-second chunks for both resting and motor imagery tasks. A clear distinction between these states by the model indicated effective learning of brain activity patterns. As shown in Fig.3(a), the cross-modal synthetic BOLD signals outperformed the real data in accuracy and precision. When combined with real data, the classification metrics improved significantly, although recall remained similar to the real data. Ablation experiments demonstrated that CAB enhances the model’s ability to use EEG features to constrain BOLD signal synthesis, improving classification metrics compared to models without the CAB. At the same time, the results in Fig.3(b) provide a direct comparison with the existing AE-based model, and the results show that the proposed CATD framework achieves superior performance across all classification metrics.
@@
Spatial (RMSE, SSIM) and temporal (cosine similarity, CCC) metrics of the synthesized BOLD signal were also calculated for the three motor imagery states. The results in Fig.4 show low RMSEs (all below 0.1) and high SSIM (all above 0.6735). The cosine similarity is approximately 0.85, and the CCC value is about 0.8, indicating a high spatio-temporal correlation with the real signal. The ablation experiments (e.g., Fig.3(a) and the No CAB section in Fig.4) confirm the effectiveness of the CAB in improving signal quality.
@@
Figure 4: Results of quantitative spatial and temporal metrics for synthetic BOLD signals in three states of motor imagery from ablation experiments. The upper half shows spatial metrics, and the lower half shows temporal metrics, with light green indicating use of CAB and teal indicating no CAB.
@@
To further illustrate the capability of the proposed model in learning high-dimensional heterogeneous brain activity features and achieving cross-modal transitions, t-SNE plots were used (Fig.5). The distribution of BOLD signals synthesized with the full CATD framework more closely matches the real signal distribution, underscoring the critical role of the CAB in enhancing signal quality. As shown in Fig.5(c), the comparison with the AE-based model demonstrates that the CATD framework not only generates a more accurate distribution of BOLD signals but also effectively captures the underlying structure of the real data, whereas the baseline method exhibits significant deviations from the real signal distribution. This further validates the advantage of the proposed approach in preserving the complex spatio-temporal relationships inherent in brain activity signals.
@@
Figure 5: (a) t-SNE plots of synthesized versus real BOLD signal distributions in the CATD framework across two motor imagery states. (b) t-SNE plot of generated versus real BOLD signal distributions after ablation of the CAB. (c) t-SNE plots of BOLD signal distributions generated by the comparison method versus real signals.
@@ 3.3 Evaluation of temporal resolution enhanced BOLD signals
To verify that the CATD framework can leverage the high temporal resolution of EEG signals to achieve the temporal super-resolution of BOLD signals, temporal resolution enhancement experiments were conducted. The proposed DTFS was used for EEG to achieve triple temporal super-resolution, meaning that the temporal resolution of the generated BOLD signal was three times that of the actual BOLD signal. Since the application scenario for temporal super-resolution typically involves enhancing the existing BOLD signal rather than generating it in the absence of a BOLD signal, the original low temporal resolution BOLD signal was used as a known condition. The enhanced high temporal resolution BOLD signal was obtained by constraining the generated signal using the low-resolution original BOLD signal. The high temporal resolution results were similarly evaluated in three different motor imagery states, as shown in Table 1.
@@
Table 1: Comparison of Time Series Similarity and SNR Between Generated and Real BOLD Signals Across Three Motor Imagery States
Time Series Similarity SNR
Cosine Similarity CCC Real Synthetic
eegNF 0.98998±0.00008 0.98589±0.00003 12.9640 12.9995±0.0103
fmriNF 0.99187±0.00007 0.98714±0.00006 13.4383 13.3112±0.0137
eegfmriNF 0.98740±0.00005 0.98209±0.00002 12.4715 13.0708±0.0533
@@
An important advantage of high temporal resolution is the potential for signal-to-noise ratio (SNR) enhancement. A higher SNR implies a greater proportion of useful information relative to noise, enabling signals with higher SNR to more reliably reflect actual physiological activity. In the experiments, the SNR values of real and generated BOLD signals were shown in Table 1. From the results, it could be seen that the signal-to-noise ratios in the two states were improved except for fmriNF, the motor imagery state. This was also consistent with the results demonstrated by t-SNE above, i.e., the results were worse in the fmriNF state and better in the other two states.
@@
Figure 6: Visualization of low-resolution real BOLD signal points and high-resolution synthetic BOLD signal curves, demonstrating consistent trends.
@@
Data visualization was also performed. In the parietal region, where the correlation of motor imagery was strong, a node at the cortex corresponding to the C3 electrode portion of the international 10-20 lead system was selected. Curve graphs were used to display the signal intensities of the generated high temporal-resolution time-series signals and the real signals at this node at the corresponding time points. As shown in Fig.6, the generated high temporal resolution signal and the real signal exhibited a high degree of similarity in trend. This further suggests that the high temporal resolution results obtained using the CATD framework can effectively reflect the trend of blood oxygenation.
3.4 Evaluation of the effect of temporal features on the generated signals
@@
Different frequency bands of EEG data reflect various types of brain activity, and these differences can significantly impact the results. The effect of EEG frequency bands on BOLD signal generated results was investigated. The EEG signals of different frequency bands were obtained by band-pass filtering, and BOLD signal generation was performed using these signals. As shown in Fig.3(b), the beta and gamma bands, which are related to motor imagery, performed better in the classification metrics. Additionally, the temporal metrics of the generated results were calculated. As shown in Fig.7, the gamma band performed well in temporal metrics, while the beta band exhibited lower temporal performance.
@@
The reasons for these results are analysed below. The brain’s state changes in preparation for or during imagined movements are generally slower and smoother, leading to event-related desynchronization (ERD) in the beta band[50]. This phenomenon may explain why the beta band excels in motor imagery and resting state categorization, even surpassing the full band’s classification performance. However, this characteristic also causes the generated results of the beta band to differ considerably in temporal similarity, resulting in a lower temporal index compared to the blood oxygenation activity reflected by the full-band brain electrical activity, i.e., the real BOLD signal. On the other hand, the gamma band is closely associated with higher cognitive and perceptual functions. In motor imagery tasks, the brain’s requirement to understand instructions and make judgments involves higher cognitive and perceptual functions [51]. Therefore, the gamma band most closely matches the real results in both the classification index and timing index. The other three low-frequency bands: delta, theta, and alpha, performed poorly on both the classification metric (e.g., Fig.3(b)) and the temporal similarity metric (e.g., Fig.7), suggesting that these bands negatively affect the generated results by containing less information about the correlation between the EEG and the BOLD signal.
@@ 3.5 Support for medical decision-making
To demonstrate the potential application of the proposed CATD framework in the medical domain, cross-modal generation experiments were conducted on Parkinson’s patients to support medical decision-making. In this part of the experiment, the framework was first trained on the NODDI dataset containing paired EEG and BOLD fMRI data from healthy subjects at resting state. Subsequently, a dataset containing only resting-state EEG data from Parkinson’s patients was used to reconstruct their cortical BOLD functional maps using the trained cross-modal generation model.
Predictions of Parkinson’s disease were performed using the recorded EEG signal, the generated BOLD signal, and their combination. The classification metrics are shown in Fig.3(c). Results indicate that the simultaneous use of generated BOLD signals and recorded EEG signals significantly improves the accuracy of Parkinson’s disease prediction and related metrics. This suggests that the CATD framework effectively captures the potential connection between BOLD signals and EEG signals and can be applied across different datasets, which is crucial for disease diagnosis. Consequently, the framework not only enhances the accuracy of existing diagnostic methods but also provides new tools and methodologies for the diagnosis and research of other neurological diseases.
In order to fully utilize the advantage of our CATD framework, i.e., to obtain BOLD signals with high spatial resolution without the condition of fMRI detection, we performed a difference analysis of generated BOLD functional maps in two healthy subjects and two Parkinson’s patients. As shown in Fig.8, both showed significant abnormalities in the brain regions marked by red circles. This region is the lingual gyrus, the medial occipito-temporal gyrus and the lateral occipito-temporal gyrus[52], which, in patients with Parkinson’s disease, shows significant structural changes[53], and in patients with PD accompanied by visual hallucinations, the atrophy of these three brain regions correlates with the severity of visual hallucinations[54]. This result demonstrates the potential application of the proposed method in localizing regions of abnormal brain activity and further proves its practical value in medical diagnosis.
@@ 4 Discussion
Cross-modal neuroimage synthesis is becoming crucial in neuroanalysis research. For the first time, we propose the CATD framework based on diffusion models to achieve cross-modal synthesis of temporal functional neuroimages, enabling the conversion of EEG to BOLD signals. This approach addresses limitations in BOLD acquisition, such as the inability of patients with metal implants to undergo fMRI scans. Enhancing BOLD temporal resolution facilitates the study of sub-millimeter cortical structures and activities[55], improves the quality and reliability of functional connectivity and task-driven fMRI research through higher sampling rates[56], and allows for more accurate detection and localization of functional activation regions, capturing transient neural activities[57]. This work provides novel insights for functional neuroimage synthesis[58]. In the following, we will analyze the experimental results in detail to illustrate the performance of our framework.
@@
In the experiments on the motor imagery dataset, our CATD framework demonstrates its superior performance in cross-modal feature learning and synthesis by detecting categorical metrics for both motor imagery states and resting states. Specifically, the framework is able to capture the features of EEG and BOLD signals well and perform effective cross-modal synthesis. This improved performance can be attributed to the ability to capture complex modes and integrate different modal features in the CATD framework, which enables it to achieve accurate transitions between various signal features. Ablation experiments further validate the effectiveness of the proposed CAB in aligning high-dimensional mismatched data pairs. CAB can learn and capture valid links between EEG and BOLD signals, ensuring the robustness of the model in cross-modal feature capture and synthesis. Compared to the AE-based model, the CATD framework achieves superior results due to the ability to extract advanced representations of brain activity and ensure better alignment of temporal and spatial features.
@@
In the temporal super-resolution experiments, the results exhibit a high correlation in all three motor imagery states, while the signal-to-noise ratio is improved in two states compared to the original signal. This suggests that temporal resolution enhancement is indeed feasible and can provide a more detailed and accurate temporal representation of the BOLD signal.
@@
In order to further verify which EEG frequency bands have a greater impact on the results and have the potential to improve the efficiency of cross-modal synthesis, a frequency band analysis was performed. It was found that the beta band had the greatest impact on the classification results, while the gamma band performed the best in terms of similarity to the real signal. This can be explained by the properties of EEG signals. Lower frequency bands such as beta show the event-related desynchronization (ERD) phenomenon in motor imagery tasks and therefore perform better in classification metrics. Whereas the higher cognitive functions involved in the motor imagery task are mainly determined by the gamma band, the generated results of the gamma band therefore show the highest similarity to the original signal.
@@
The CATD framework’s ability to synthesize BOLD signals from EEG has significant implications for clinical practice. By providing a non-invasive and cost-effective alternative to conventional fMRI, the framework can increase accessibility for patients who cannot undergo fMRI scans, such as those with metal implants. At the same time, the improved temporal resolution and functional maps generated by the framework can facilitate early diagnosis and intervention in neurological disorders like epilepsy and Parkinson’s disease. For example, identifying abnormal brain activity regions with high temporal and spatial resolution could help refine therapeutic strategies. The ability to monitor disease progression using synthetic neuroimages also offers potential for personalized medicine approaches.
An exciting direction for future research is the potential to extend the CATD framework to predict neural activity. Leveraging EEG signals and integrating temporal forecasting techniques within the CATD framework could enable the anticipation of upcoming BOLD signal patterns. While the current implementation does not achieve real-time generation due to the computational complexity of diffusion-based models, advancements in GPU and hardware acceleration technologies could enable real-time prediction in the near future. Real-time capabilities would significantly enhance the framework’s applicability in dynamic brain monitoring, real-time neurofeedback systems, and adaptive brain-machine interfaces.
While the results are promising, achieving accurate mapping from EEG to BOLD signals still presents challenges. Fig.9 illustrates a typical failure case, where the lower regions of the generated BOLD map significantly differ from the real image. This discrepancy is likely due to the low spatial resolution and relatively low signal-to-noise ratio (SNR) of EEG signals. Even though our model successfully learns the relationship between EEG and BOLD signals, the process of upsampling to generate high-resolution images introduces inaccuracies. Addressing these issues of spatial resolution and SNR will be crucial for enhancing the precision of cross-modal synthesis in future work.
Figure 9: Typical failure case of our method, showing significant differences in the lower BOLD functional map (near the frontal lobe).
In addition to the failure case highlighted in Fig.9, other limitations exist. First, the limited number of paired data subjects constrained the training process. To address the limitations of dataset size, we have carefully selected, to the best of our knowledge, the relatively larger datasets among the publicly available ones with simultaneously acquired EEG and fMRI signals. However, the limited number of paired data subjects remains a constraint and poses challenges to achieving even greater robustness and generalizability. Introducing pre-trained models into the CATD framework could address the data scarcity issue by leveraging transfer learning to improve representation accuracy and enhance the model’s feature extraction capability. Second, inter-subject variability in EEG and BOLD signal patterns poses a challenge for achieving consistent cross-modal synthesis across diverse populations. Finally, the computational complexity of the CATD framework, particularly with its reliance on the CAB module and diffusion model, may hinder its scalability for large-scale clinical applications. Addressing these challenges in future work will be essential to fully realize the framework’s potential in both research and clinical settings.